1. How many combination are there of 9 identical cars if I am picking 3 of them to test drive?
2. A garden has a total of 10 flowers. Three are yellow, two are white, and five are blue. If you select 2 at random without replacement, find the probability that
a. non of the flowers are yellow
b. at least one of the flowers are yellow

Answer by stanbon(75887)

1. How many combination are there of 9 identical cars if I am picking 3 of them to test drive?
9C3 = (9*8*7)/(1*2*3) = 84
----------------------------------
2. A garden has a total of 10 flowers. Three are yellow, two are white, and five are blue. If you select 2 at random without replacement, find the probability that
a. none of the flowers are yellow::: 7/10
-------------------------------------------------
b. at least one of the flowers is yellow
P(at least one) = 1 - P(none is yellow)
= 1 - 0.7
= 0.3
====================
Cheers,
Stan H.
==========